# PAT: Pixel-wise Adaptive Training for Long-tailed Segmentation

### v2

@@ Abstract
Beyond class frequency, we recognize the impact of class-wise relationships among various class-specific predictions and the imbalance in label masks on long-tailed segmentation learning. To address these challenges, we propose an innovative Pixel-wise Adaptive Training (PAT) technique tailored for long-tailed segmentation. PAT has two key features: 1) class-wise gradient magnitude homogenization, and 2) pixel-wise class-specific loss adaptation (PCLA). First, the class-wise gradient magnitude homogenization helps alleviate the imbalance among label masks by ensuring equal consideration of the class-wise impact on model updates. Second, PCLA tackles the detrimental impact of both rare classes within the long-tailed distribution and inaccurate predictions from previous training stages by encouraging learning classes with low prediction confidence and guarding against forgetting classes with high confidence. This combined approach fosters robust learning while preventing the model from forgetting previously learned knowledge. PAT exhibits significant performance improvements, surpassing the current state-of-the-art by $2.2\%$ in the NyU dataset. Moreover, it enhances overall pixel-wise accuracy by $2.85\%$ and intersection over union value by $2.07\%$, with a particularly notable declination of $0.39\%$ in detecting rare classes compared to Balance Logits Variation, as demonstrated on the three popular datasets, i.e., OxfordPetIII, CityScape, and NYU. The code will be released upon acceptance.
@@
Long-tailed Learning Image Segmentation Adaptive Training
@@ 1 Introduction
The integration of deep learning (DL) into real-world applications has sparked fresh enthusiasm for research extending beyond meticulously crafted datasets with balanced classes and accurate representations of the test set distribution. However, it is doubtful to foresee and construct datasets that encompass all potential scenarios consistently. Therefore, it is crucial to explore robust algorithms that can perform well on imbalanced datasets and unforeseen circumstances during testing. These challenges can be classified as distributional shifts between the training and testing conditions, notably, label prior shift and non-semantic likelihood shift [8].
Many researchers attempt to propose robust algorithms against long-tailed rare categories in segmentation via resampling [3, 7], data augmentation [26, 20], logits adjustment (LA) [9, 1, 24, 12, 27], domain adaptation (DA) [30, 11, 29]. However, the imbalance among classes inside samples in segmentation remains a critical issue. Existing efforts predominantly concentrate on addressing sample imbalance within classes, with limited research devoted to long-tailed segmentation. The data augmentation and sampling method prove inadequate as it can neither augment nor sample classes within masks. Additionally, designing model architectures for this purpose is highly resource-intensive and computationally demanding.
@@
Recognizing these challenges, we delve into existing research on long-tailed segmentation and uncover intriguing insights. 1) Imbalanced mask representations: beyond the difficulties posed by rare objects, imbalanced mask representations occur when some masks dominate the learning process, leading to a bias towards recognizing dominant classes and neglecting minority classes. 2) Model uncertainty and degradation: models facing uncertainty often produce low-precision channel-wise logits, leading to biased gradient updates. These updates favor incorrect label predictions and ignore progress toward the true labels, further degrading performance.
Building upon these insights, we introduce Pixel-wise Adaptive Training (PAT), a novel approach for addressing long-tailed rare category problems in segmentation. PAT comprises two key contributions: 1) Class-wise Gradient Magnitude Homogenization: We address the imbalanced learning caused by differences in object size across classes by dividing the loss by the corresponding class mask’s size. This effectively equalizes the influence of each class on the learning process. 2) Pixel-wise Class-Specific Loss Adaptation (PCLA): This component focuses on pixel-wise predicting vectors (PPVs) within each pixel (see Fig. 2(a)). By examining the PPVs, we can evaluate how individual channels influence the learning process by analyzing the logit predictions. Specifically, by employing an inverted softmax function, we can strike a balance between two factors: the presence of long-tailed rare objects and the impact of insufficient loss contribution on the joint loss function. This equilibrium enables us to determine coefficients that prioritize learning in classes with rare objects or those with minimal contributions to the joint loss induced by the wrong predictions of the model from previous low-performance learning progress. Consequently, we address the issue of imbalanced learning stemming from both current training samples and previously learned imbalances in the models.
@@ 2 Related Works
2.1 LongTailed Learning in Segmentation
Several studies have explored the phenomenon of longtail distribution within segmentation contexts. Different from the traditional classification task, resampling methods [3, 7] as well as data augmentation [26, 20] cannot tackle the underlying issues of long-tail distribution in segmentation as the ratio among class frequencies is not adjusted[31]. Logits adjustment (LA) is one of the most prominent techniques that can balance the effect of head class logits by equalizing the output logits distribution [9, 1, 24, 12, 27]. Domain adaptation [30, 11, 29] is also considered to enhance the logits balancing, though accessing the target dataset is impractical in real-world applications [25]. Post-processing is useful in removing the uncertainty [18] at the pixel level in semantic segmentation, though it requires a calibration process for each dataset.
2.2 Class Sensitive Learning
Class-sensitive learning (CSL) is one direction in solving long-tailed learning [31], which takes full advantage of classes’ frequency to balance the logits distribution, class-wise gradient, etc. From Section 2.1, it turns out that LA and CSL are the two most easy-to-use yet effective methods in long-tailed learning. Focal (Focal) function [13] is firstly proposed to tackle this challenge by taking the inverse of logits as a weight for each class. In [6], a class-balance loss function is proposed, working as an effective sampler by a class frequency weighting function. The combination of class-balance loss and Focal is also considered in [6]. Balancing the effect of the exponential function in the Softmax activation function is studied [24, 12]. Other recent methods [4, 26] focus on using noise in logits to balance them which are also popular in segmentation, we also compare our proposed method with those ones. We also show that our proposed method surpasses the others in both performance and utilization.
@@ 3 Proposed Method
Figure 1: Overall methodology. 1) Training procedure: an input image $x_{i}$ is fed into a typical encoder-decoder model architecture to produce output $\hat{x}_{i}$. This prediction’s logits are then adjusted to create a weight tensor, which has the same size as $\hat{x}_{i}$. The normalized $\hat{x}_{i}$ is multiplied by the weight tensor to balance the dominant logits. Finally, the loss value is obtained using the proposed PAT loss function. 2) Logits Adjustment: The logits vector is normalized by the Softmax function, added by a tensor of $-1$, and scaled by the exponential function to find the inverse dominant coefficients $\beta_{i,j}$. Then, these coefficients are normalized into a range of $[0,1]$ to form the weight tensor.
@@ 3.1 Problem Statement
Consider $(x,y)\sim P(\mathcal{X},\mathcal{Y})$, where $x\in\mathbb{R}^{N\times C\times H\times W}$ and $y\in\mathbb{R}^{N\times L\times H\times W}$ are input data and corresponding ground truth, respectively. $C$, $L$ are the number of image channels and categories, respectively. $N$, $H$, and $W$ are the total number of training samples, height, and width of the image, separately. The segmentation problem is represented in Eq. (1).
@@ 3.2 Imbalance among label masks
One major challenge in image segmentation is the class imbalance in label masks (see Fig. 2). Larger masks contribute more significantly to the loss of function than smaller masks, leading to a bias towards dominant classes. Specifically, we come over the class-wise loss component, which can be represented as:
@@
Figure 2: Quantitative analysis on the imbalance in mask size among classes. The vertical axis illustrates the mask size calculated by the total number of pixels, which are associated with the corresponding mask. The horizontal axis shows different masks that potentially appear in the ground truth. 1(a)) While road and vegetation take 50000 pixels and 60000 pixels, respectively, cars account for around 1000 pixels. 1(b)) Not only road and vegetation but also walking and sky take roughly $99\%$ proportion, compared to cars. 1(c)) Void and build take nearly 70000 pixels and 60000 pixels, under 1000 pixels are accounted by sign, and no car appears. 1(d)) Cars take roughly 20000 pixels, though their masks’ sizes are much bigger than the ones in Figs. 1(a) and 1(b).
@@ 3.3 Pixel-wise Adaptive Traning with Loss Scaling
where $\hat{x}^{l}_{i,j}$ is the logits prediction of sample $i$ at pixel $j$ with regard to category $l$. In segmentation problems, the class is considered by a composition of many pixels over the masks. We hypothesize that the learning in each class may occur diversely according to the classification of different pixels. Therefore, we propose the pixel-wise adaptive (PAT) loss via the pixel-wise adaptive coefficient set $\beta_{i,j}\in\mathbb{R}^{L}=\large[\beta_{i,j}^{0},\beta_{i,j}^{1},\cdots,\beta_{i,j}^{L-1}\large]$.
$\mathcal{L}(x,y)=-\frac{1}{B}\sum^{B-1}_{i=0}\sum^{HW}_{j=1}\Biggr{[}\sum^{L-1}_{l=0}\beta_{i,j}^{l}\times\frac{y^{l}_{i,j}}{S^{l}_{i}}\log\Big{(}\frac{\exp\{\hat{x}^{l}_{i,j}\}}{\sum_{l^{\prime}=0}^{L-1}\exp\{\hat{x}^{l^{\prime}}_{i,j}\}}\Big{)}\Biggr{]}.$ (4)
@@
where $p(\hat{x}_{i,j})=\large[\exp\{\hat{x}^{l}_{i,j}\}/\sum_{l^{\prime}=0}^{L-1}\exp\{\hat{x}^{l^{\prime}}_{i,j}\}\large]_{l=0}^{L-1}$ indicates the output of the softmax activation which normalize the logits vector elements into probability range of $[0,1)$. We define $T$ as a temperature coefficient and $\epsilon$ as an arbitrary constant. By tuning the $\beta_{i,j}$ according to each pixel-wise vector $\hat{x}_{i,j}=\{\hat{x}^{l}_{i,j}|~{}l\in\{0,\ldots,L-1\}\}$, our proposed coefficient hinges on two key concepts: firstly, equalizing the loss value across various logits, and secondly, preventing negative transfer in well-classified results. Further analysis is presented in Section 4.2. Additionally, in Eq. (4), we normalize the loss across all components by dividing by $S^{l}_{i}$. This approach allows us to penalize loss components that have a dominant size relative to others, thereby facilitating the homogenization of class-wise gradient magnitudes.
4 Theoretical Analysis
@@ 4.2 Analysis on the PAT to the logits imbalance
To have a comprehensive understanding of PAT robustness to the imbalance rare object segmentations, we compare the loss value of PAT and Focal[13] at different logit probabilities in Fig. 2(c), yielding two significant observations. Firstly, by parameterizing the loss function with the PAT scaling factor, we observe more balanced learning across classes with varying logit probabilities. Secondly, in contrast to Focal, we notice that losses associated with high-probability logits are not nullified. This phenomenon can prevent negative transfer on well-classified samples. In comparison, the PAT-scaling parameterized loss function fosters equitable learning while retaining loss information from high-probability logits, thus allowing the training process to retain valuable knowledge from well-classified samples. Consequently, PAT learning is anticipated to exhibit greater stability compared to Focal techniques.
4.3 Analysis on the PAT to the gradient magnitude homogenization
Magnitude differences of the gradients across tasks may lead to a subset of tasks dominating the total gradient, and therefore to the model prioritizing them over the others [10]. This phenomenon happens obviously in the segmentation task. To have a comprehensive understanding, from Eq. (2) we consider the gradient across classes as follows:
$\nabla\mathcal{L}_{l}(x,y)=S^{l}_{i}\cdot\nabla\bar{\ell}_{l}(x_{i},y_{i}).$ (6)
Therefore, the gradient norm proportion between classes $l$ and $l^{\prime}$ is as follows:
$\frac{\|\nabla\mathcal{L}_{l}(x,y)\|}{\|\nabla\mathcal{L}_{l^{\prime}}(x,y)\|}=\frac{S^{l}_{i}\cdot\|\nabla\bar{\ell}_{l}(x_{i},y_{i})\|}{S^{l^{\prime}}_{i}\cdot\|\nabla\bar{\ell}_{l^{\prime}}(x_{i},y_{i})\|}$ (7)
As the mask size $S^{l}_{i}$ becomes divergent across classes in one task, the gradient magnitude is diverse, thereby biasing the learning towards the subset of classes with dominant mask sizes. By applying the normalization via mask size $S^{l}_{i}$, as mentioned in Eq. (4), we can have the model update focus on the labels’ canonical loss ¯$\bar{\ell}_{l}(x_{i},y_{i})$. This approach bears resemblance to gradient magnitude normalization techniques discussed in [10, 15], facilitating more balanced contributions among classes during learning. Nevertheless, by normalizing directly according to the mask size, we can notably decrease computational complexity compared to utilizing gradient norms, as seen in the previously mentioned studies.

### v4

@@ Abstract
Beyond class frequency, the impact of class-wise relationships among various class-specific predictions and the imbalance in label masks can cause significant problems in long-tailed segmentation learning. Addressing these challenges, we propose an innovative Pixel-wise Adaptive Training (PAT) technique tailored for long-tailed segmentation. PAT has two key features: 1) pixel-wise class-specific loss adaptation (PCLA), 2) head and tail balancing, and 3) low computation cost. First, PCLA tackles the detrimental impact of both rare classes within the long-tailed distribution and inaccurate predictions from previous training stages by encouraging learning classes with low prediction confidence. Second, PAT integrates a new weighting curve function for guarding against forgetting classes with high confidence. Third, PAT takes advantage of a pixel-wise weighting mechanism thus requiring computation cost just above Cross-Entropy. PAT exhibits significant performance improvements, surpassing the current state-of-the-art by $2.2\%$ in the NyU dataset. Moreover, it enhances overall pixel-wise accuracy by $2.85\%$ and intersection over union value by $2.07\%$, with a particularly notable declination of $0.39\%$ in detecting rare classes compared to Balance Logits Variation, as demonstrated on the three popular datasets, i.e., OxfordPetIII, CityScape, and NyU. The code is available at https://github.com/KhoiDOO/ibla.
@@
Deep Learning , Segmentation , Long-tailed Learning , Loss Function , Class Sensitive Learning
journal: Pattern Recognition Letters
[…]
{highlights}
Long-tailed rare objects cause unstable training Segmentation models.
Long-tailed datasets consist of imbalanced frequencies among masks and inside masks.
Adaptive weight from Class-sensitive learning loss function balance gradient learning.
Putting more weight on small objects while not forgetting high-confidence objects.
Low computation cost loss function for large-scale models and datasets.
@@ 1 Introduction
Figure 1: Quantitative analysis on the imbalance in mask size among classes. The vertical axis illustrates the mask size calculated by the total number of pixels. The horizontal axis shows different masks that potentially appear in the ground truth. 1(a)) While road and vegetation masks take 50000 pixels and 60000 pixels, respectively, cars account for around 1000 pixels.
The applications of deep learning (DL) have shifted research interest toward datasets that are not perfectly balanced or meticulously crafted. It is crucial to explore robust algorithms that can perform well on imbalanced datasets. These challenges can be classified as distributional shifts between the training and testing sets[1]. Resampling [2, 3], data augmentation [4, 5], logits adjustment (LA) [6, 7, 8, 9, 10], and domain adaptation (DA) [11, 12, 13] have been used to address long-tailed rare categories in segmentation. However, existing methods primarily focus on sample imbalance within classes, overlooking the issue of class imbalance within segmentation samples.
@@
The drawbacks above originated from the following insights. 1) Imbalanced mask representations: beyond the difficulties posed by rare objects, imbalanced mask representations occur when some masks dominate the learning process, leading to a bias towards recognizing dominant classes[14, 4, 8]. 2) Model uncertainty and degradation: models facing uncertainty often produce low-precision channel-wise logits, leading to biased gradient updates. These updates favor incorrect label predictions and ignore progress toward the true labels, further degrading performance[15, 16, 17, 18]. 3) High confidence categories neglection: Current approaches focus on putting a higher weight on long-tailed rare objects while putting strictly small or zeroing weight causing unstable learning[15]. 4) High computation cost: Although current approaches can tackle the issues[17, 4, 8], they cost a large amount of computation.
We introduce Pixel-wise Adaptive Training (PAT), a novel approach for addressing long-tailed rare categories in segmentation. PAT comprises the following key contributions: 1) Pixel-wise Class-Specific Loss Adaptation (PCLA): This component focuses on pixel-wise predicting vectors (PPVs) within each pixel (see Fig. 3(a)). By examining the PPVs, we can evaluate how individual channels influence learning by adjusting the logit predictions. 2) Head and Tail Balancing: By employing an inverted softmax function, we can balance learning the presence of long-tailed rare objects and maintaining the model performance on high-confidence categories. 3) Low computation cost: By taking full advantage of the Focal[15] weighting mechanism, we propose a new weighting curve that figures out long-tailed learning while requiring no supporting tensors, resulting in an extremely lower computation cost.
@@ 2 Related Works
Different from the traditional classification task, resampling methods[2, 3] and data augmentation[4, 5, 19] cannot tackle the issues of long-tail distribution in segmentation as the ratio among class frequencies is not adjusted[14].
Multi-stage & Multi Objective Optimization. Multi-stage solution[20, 21, 22, 23] consists of training feature extractor to learn representation of imbalance datasets, and fine-tuning classifier via frozen feature extractor. Otherwise, multi-objective optimization[24, 25, 26, 27], combining cross-entropy loss with contrastive loss or regularization.
Logits adjustment. Logits adjustment (LA) is one of the most prominent techniques that can balance the effect of head class logits by equalizing the output logits distribution[6, 7, 8, 9, 10]. Domain adaptation[11, 12, 13] is also considered to enhance the logits balancing, though accessing the target dataset is impractical in real-world applications[28]. Post-processing is useful in removing the uncertainty[29] at the pixel level in semantic segmentation, though it requires a calibration process for each dataset.
Class Sensitive Learning. Class-sensitive learning (CSL) is one direction in solving long-tailed learning[14], which takes full advantage of classes’ frequency to balance the logits distribution, class-wise gradient, etc. LA and CSL are the two most easy-to-use yet effective methods in long-tailed learning. Focal (Focal) function[15] is firstly proposed to tackle this challenge by taking the inverse of logits as a weight for each class. In[16], a class-balance loss function is proposed, working as an effective sampler by a class frequency weighting function. The combination of class-balance loss and Focal is also considered in[16]. Balancing the effect of the exponential function in the Softmax activation function is studied[8, 9]. Other recent methods[17, 4] focus on using noise in logits to balance them which are also popular in segmentation, we also compare our proposed method with those. We also show that our proposed method surpasses the others in both performance and utilization.
@@ 3 Proposed Method
Figure 2: Overall methodology. 1) Training procedure: an image $x_{i}$ is fed into an encoder-decoder architecture to produce logits $\hat{x}_{i}$. Subsequently, $\hat{x}_{i}$ is adjusted to create a weight tensor, which has the same size as $\hat{x}_{i}$. The normalized $\hat{x}_{i}$ is multiplied by the weight tensor to equalize the contribution of each logit to the PAT loss function. 2) Logits Adjustment: Logits vector is normalized by the Softmax, added by a tensor of $-1$, and scaled by the exponential function to find the inverse dominant coefficients $\beta_{i,j}$. Then, $\beta_{i,j}$ is normalized to $[0,1]$ to form the weight tensor.
@@ 3.1 Problem Statement
Consider $(x,y)\sim P(\mathcal{X},\mathcal{Y})$, where $x\in\mathbb{R}^{N\times C\times H\times W}$ and $y\in\mathbb{R}^{N\times L\times H\times W}$ are input data and corresponding ground truth, respectively. $C$ and $L$ are the number of image channels and categories, respectively. $N$, $H$, and $W$ are the total number of training samples, height, and width of the image, separately. The segmentation problem is represented in Eq. (1).
@@ 3.2 Imbalance among label masks
One major challenge in image segmentation is the class imbalance in label masks. Larger masks contribute more significantly to the loss of function than smaller masks, leading to a bias towards dominant classes. Specifically, we come over the class-wise loss component, which can be represented as:
@@ 3.3 Pixel-wise Adaptive Traning with Loss Scaling
where $\hat{x}^{l}_{i,j}$ is the logits prediction of sample $i$ at pixel $j$ with regard to category $l$. In segmentation problems, the class is considered by a composition of many pixels over the masks. We hypothesize that the learning in each class may occur diversely according to the classification of different pixels. Therefore, we propose the pixel-wise adaptive (PAT) loss via the pixel-wise adaptive coefficient set $\beta_{i,j}\in\mathbb{R}^{L}=[\beta_{i,j}^{0},\beta_{i,j}^{1},\cdots,\beta_{i,j}^{L-1}]$.
$\mathcal{L}(x,y)=-\frac{1}{B}\sum^{B-1}_{i=0}\sum^{HW}_{j=1}\sum^{L-1}_{l=0}\beta_{i,j}^{l}\times\frac{y^{l}_{i,j}}{S^{l}_{i}}\log\Big{(}\frac{\exp\{\hat{x}^{l}_{i,j}\}}{\sum_{l^{\prime}=0}^{L-1}\exp\{\hat{x}^{l^{\prime}}_{i,j}\}}\Big{)}.$ (4)
@@
where $p(\hat{x}_{i,j})=\left[\exp\{\hat{x}^{l}_{i,j}\}/\sum_{l^{\prime}=0}^{L-1}\exp\{\hat{x}^{l^{\prime}}_{i,j}\}\right]_{l=0}^{L-1}$ indicates the output of the softmax activation which normalize the logits vector elements into probability range of $[0,1)$. We define $T$ as a temperature coefficient and $\epsilon$ as an arbitrary constant. By tuning the $\beta_{i,j}$ according to each pixel-wise vector $\hat{x}_{i,j}=\{\hat{x}^{l}_{i,j}|~{}l\in\{0,\ldots,L-1\}\}$, our proposed coefficient hinges on two key concepts: firstly, equalizing the loss value across various logits, and secondly, preventing negative transfer in well-classified results. Further analysis is presented in Section 4.2. Additionally, in Eq. (4), we normalize the loss across all components by dividing by $S^{l}_{i}$. This approach allows us to penalize loss components that have a dominant size relative to others, thereby facilitating the class-wise gradient magnitudes.
4 Empirical Analysis
In Section 4.1, we demonstrate that our proposed PAT effectively addresses the numerical instability that may arise due to Eq. (5). In Section 4.2, we prove that PAT can effectively handle long-tailed rare object segmentation, particularly in detecting objects with small portions of the mask. Moreover, PAT maintains the performance of high-confidence classes. In Section 4.3, we show that PAT achieves low computational costs compared to existing state-of-the-art methods by incorporating a simple weighting mechanism that does not require additional supporting tensors or calculations.
@@ 4.2 Analysis on the PAT to the logits imbalance
Experimentally, Focal loss performance is lower than current approaches, through its simple and optimized implementation. To have a comprehensive understanding of PAT robustness to the imbalance rare object segmentations, we compare the loss value of PAT and Focal[15] at different logit probabilities in Fig. 3(c), yielding two significant observations.
Table 1: Adaptive loss value comparison between Focal loss and PAT loss functions with different adaptive coefficients: $\gamma\in\{2,5\}$ and $T\in\{2,5\}$.
$\boldsymbol{p(\hat{x}_{i,j})}$ 0.2 0.3 0.4 0.5
Focal ($\gamma=2$) 1.03 0.59 0.33 0.17
PAT ($T=2$) 1.08 0.85 0.68 0.54
Focal ($\gamma=5$) 0.53 0.2 0.07 0.02
PAT ($T=5$) 1.37 1.05 0.81 0.63
$\boldsymbol{p(\hat{x}_{i,j})}$ 0.6 0.7 0.8 0.9
Focal ($\gamma=2$) 0.08 0.03 0.01 0.0
PAT ($T=2$) 0.42 0.31 0.2 0.1
Focal ($\gamma=5$) 0.01 0.0 0.0 0.0
PAT ($T=5$) 0.47 0.34 0.21 0.1
Firstly, by parameterizing the loss function with the PAT scaling factor, we observe more balanced learning across classes with varying logit probabilities. Secondly, in contrast to Focal, we notice that losses associated with high-probability logits are zero-weighted (refers to Tab. 1). This phenomenon can prevent positive transfer on well-classified samples. In comparison, the PAT-scaling parameterized loss function fosters equitable learning while maintaining loss information for high-probability logits, thus enabling stable training.
4.3 Analysis on time and space complexity
As demand for low-cost computation algorithm[14, 30, 31], we take into consideration time and space complexity. Tab. 2 theoretically suggests that traditional Cross-Entropy, Focal, and PAT have the lowest complexity in both time and space, compared to LDAM and BLV, which require extensive supporting tensors ($\Delta_{y}$, $\delta(\sigma)$, and $c$) to manipulate margin probability distribution.
Table 2: Theoretical time ($\mathcal{O}$) and space ($\mathcal{V}$) complexity comparison between state-of-the-arts and PAT. Denote $\mathcal{F}$ as loss formula, $\gamma$ is scale coefficient of Focal[15], $z_{y}$, $z_{i}$ are logits, $\sigma$ is standard deviation, $\delta$ is a distribution generator, $\psi=BCHW$, and $\bar{p}(x)=1-p(x)$
Method $\boldsymbol{\mathcal{F}}$ $\boldsymbol{\mathcal{O}}$ $\boldsymbol{\mathcal{V}}$
CE $-\log(p(x))$ $\mathcal{O}(\psi)$ $\mathcal{O}(\psi)$
Focal $-(\bar{p}(x))^{\gamma}\log(p(x))$ $\mathcal{O}(\psi)$ $\mathcal{O}(\psi)$
LDAM $-\log(\frac{\exp(z_{y}-\Delta_{y})}{\sum_{i}\exp(z_{i}-\Delta_{i})})$ $\mathcal{O}(2\psi)$ $\mathcal{O}(2\psi)+\mathcal{O}(2C)$
BLV $-\log(\frac{\exp(z_{y}+c_{y}\delta(\sigma))}{\sum_{i}\exp(z_{i}+c_{i}\delta(\sigma))})$ $\mathcal{O}(2\psi)$ $\mathcal{O}(3\psi)+\mathcal{O}(2C)$
PAT $-\exp\{(\bar{p}(x))/T\}\log(p(x))$ $\mathcal{O}(\psi)$ $\mathcal{O}(\psi)$